# High-performance superconducting two-qubit gate using a three-mode tunable coupler

### v1

@@ III.2 CZ gate concept
Since changing the external magnetic flux in the coupling element, directly controls the ZZ interaction between qubits, the idea of the controlled-phase gate implementation looks quite natural. One can switch on the interaction adiabatically increasing the external flux and, after waiting for the required time, return the coupler to the initial state. The flux pulse is supposed to be short while remaining adiabatic in order to prevent undesired excitation transitions between the computational states and leakage out of the computational subspace. Therefore, in order to activate the interaction we use a smooth pulse $\Phi_{\text{ext}}(t)$ with a Hann function-shaped edges of 9 ns both in the simulation and the experiment, as it is shown in Fig. 3(d). Also, we define the gate duration as the length of the entire pulse, accounting both edges and a plateau-shaped part.
@@
We described above in details the method of constructing the system model using two-tone spectroscopy data and measurement of the ZZ-interaction strength at the zero-flux point. Using the obtained values, we construct the full-circuit Hamiltonian and calculate the time dynamics of the system during the CZ gate implementation. In this numerical analysis, we take into account the first four energy levels of each qubit and each mode of the coupler in order to reduce a computational time. Thus, the Hilbert space dimension of the full system in the simulation is $4^{5}$. We denote the states of each computational qubit as $|g\rangle,|e\rangle,|f\rangle$ and use numerical notation for coupler modes.
@@
We simulate the CZ gate performance for an amplitude of the coupler’s SQUID flux pulse in a range from $0.37\Phi_{0}$ to $0.5\Phi_{0}$. For each amplitude value we find a gate duration, making the state $|ee\rangle$ accumulated conditional phase $\theta=\pi$. In order to evaluate the performance of the resulting CZ operation, we use the standard expression for the two-qubit gate fidelity [42, 43]:
$F=\frac{\text{Tr}(R_{\text{ideal}}^{\dagger}R)+D}{D(D+1)}.$ (3)
Here, $D=2^{n}$ is the dimension of the computational Hilbert subspace ($n=2$), $R_{\text{ideal}}$ and $R$ are Pauli transfer matrices corresponding to the action of the ideal and simulated gates, respectively. Moreover, we calculate an average leakage out of the computational subspace $L=1-L_{e}$, where $L_{e}$ is a leakage error, and the mean projection of computational states population onto the target states after the CZ operation for each four states:
$\begin{split}P_{t}=\cfrac{1}{4}\sum_{\psi}|\langle\hat{\mathrm{CZ}}(\psi)|\hat{\mathcal{U}}(\psi)\rangle|^{2}=|\langle\psi|\hat{\mathcal{U}}(\psi)\rangle|^{2},\\ \psi\in\{|gg\rangle,|ge\rangle,|eg\rangle,|ee\rangle\}.\end{split}$ (4)
Here the unitary operator $\hat{\mathcal{U}}$ denotes the time evolution of the initial computational states during the gate implementation. The result is presented in Fig. 3(a), the corresponding gate duration as a function of an external flux amplitude is given in Fig. 3(b). Here we observe a periodic behavior of the gate fidelity. At the first maximum of the fidelity curve we do not get high-fidelity operation due to the immense ZZ interaction. Withal, the second maximum, marked with a star, shows the fidelity of 0.999(7) with 43.3-ns duration. For this point, a time evolution of the mean target state population (dashed blue), a conditional phase (red) and a leakage error (dashed-dotted green) are depicted in Fig. 3(c). The target state population behavior mostly relates to the oscillations between states during the gate implementation. The population leaks to the second excited qubit states $|gf\rangle,|fg\rangle$ and to the coupler state $|010\rangle$. The detailed dynamic of each computational state can be found in Appendix B.
@@
III.3 CZ gate implementation
We were able to experimentally implement the proposed CZ gate of the total 60-ns pulse duration with $-0.28$ V amplitude. The pulse edges length is $9$ ns, that is also included in the total gate duration. Since the experimental ZZ interaction strength is larger than the approximated one, the corresponding gate evolution in the experiment is longer. Thereby, we consider that the performed gate corresponds to the point nearby the third peak in Fig. 3(a). In order to validate the two-qubit gate performance we start with the quantum process tomography experiment, the result of which is depicted in Fig. 4(c). The fidelity of the reconstructed CZ Pauli transfer matrix is $83.3\%$. We explain the reduction of the gate quality by the poor multi-qubit readout. Therefore, we utilize cross-entropy benchmarking (XEB) [1] resilient against SPAM errors in order to characterize the performance of the implemented CZ operation.
For the XEB we execute the gate sequence shown in Fig. 4(a), where $C_{ij}$ are randomly chosen single-qubit Clifford gates. Here the first index corresponds to the qubit number and the second index refers to the serial number of the gate in the sequence. The single-qubit Clifford group is constructed of the $\pi/2$ rotations around X axis of $46.66$ ns duration and virtual S gates in the same way as it was performed in Ref. [44]. First, we accomplish reference XEB consisting only single-qubit operations simultaneously implemented on both qubits. The obtained depolarization fidelity averaged over randomly sampled quantum circuits as a function of a circuit depth $m$ is shown in Fig. 4(b) with a blue color. We approximate the data with the function $ap^{m}$, where $p$ is the depolarizing parameter and $a$ is a fitting parameter used to take into account a state preparation and measurement (SPAM) errors. The resulted depolarization parameter $p_{1}$ is $0.9714\pm 0.0005$. The similar data collected for the same circuits with the interleaved CZ gate is depicted in Fig. 4(b) in green. The depolarization parameter $p_{2}$ is equal to $0.947\pm 0.001$. In order to calculate the conventional fidelity $F$ of a target two-qubit gate we use the formula [1, 45]
@@ III.3 CZ gate implementation
where $p=p_{2}/p_{1}$. The obtained CZ gate fidelity $F_{\mathrm{CZ}}$ is $(98.14\pm 0.12)\%$.
Since we use the XEB method, we can utilize it to estimate the acquired two-qubit phase $\theta$. By optimizing the linear cross-entropy between the experimental data and the simulated quantum circuits with an interleaved controlled-phase gate, we obtain $\theta=(1.03\pm 0.03)\pi$. Next, we leverage a speckle purity benchmarking to estimate the decoherence error [1]. The calculated square root of purity is $0.955\pm 0.001$. By comparing this value to $p_{2}$, it can be concluded that the major cause of the gate infidelity is a decoherence. Moreover, the CZ gate fidelity is higher than the single-qubit ones since CZ gate is shorter, see Table 1.
$f_{q}$ $T_{2}^{*}\;(\mu s)$ $T_{1}\;(\mu s)$ $t_{1Q}$ (ns)
Qubit A 6.450 8.6 15.8 $2\times 46.66$
Qubit B 6.494 10.8 8.3 $2\times 46.66$
Table 1: The device parameters: qubits frequency, dephasing and coherence times, and the length of single-qubit Clifford gates in Fig. 4(a). The coefficient 2 denotes the fact, that we construct the Clifford group out of two $R_{\mathrm{X}}(\pi/2)$ gates with duration of 46.66-ns.
@@ Acknowledgments
The authors acknowledge Alexey Ustinov for fruitful discussions and useful comments on the manuscript. The work was supported by Rosatom in the framework of the Roadmap for Quantum computing (Contract No. 151/21-503, December 21, 2021 dated 12/21/2021).
@@ Appendix A FULL-CIRCUIT HAMILTONIAN
Josephson energy (GHz) Capacitance (fF) Capacitance (fF)
$E_{J1},\;E_{J10}$ 124.2 $C_{2}$ 75.6 $C_{35}$ 0.005
$E_{J2},\;E_{J9}$ 43.5 $C_{3}$ 47.2 $C_{56}$ 9.5
$E_{J3},\;E_{J8}$ 17.4 $C_{4}$ 144 $C_{J1},\;C_{J10}$ 11.1
$E_{J4},\;E_{J7}$ 124.2 $C_{5}$ 46.5 $C_{J2},\;C_{J9}$ 3.9
$E_{J5}$ 79.5 $C_{6}$ 75.9 $C_{J3},\;C_{J8}$ 1.5
$E_{J6}$ 129.1 $C_{23}$ 9 $C_{J4},\;C_{J7}$ 11.1
- - $C_{26}$ 0.015 $C_{J5}$ 11.5
- - $C_{34},\;C_{45}$ 13 $C_{J6}$ 7.1
Table 2: Target circuit parameters of the device: Josephson energies of junctions, node and Josephson junction capacitances. The coupler Josephson junction capacitances $C_{J4}$, $C_{J5}$ and $C_{J6}$, $C_{J7}$ are taken into account in corresponding values $C_{34}$, $C_{4}$, $C_{45}$.
@@
The single-qubit Hamiltonian has the following form [38]
@@
We show the flux dependence of the coupler energy level structure obtained by numerical calculation in Fig. 6. There is an alternation of tunable and non-tunable energy levels that is due to the different contribution of the center node with the coupler SQUID.
@@
$\begin{split}T=\frac{\Phi^{2}_{0}}{2(2\pi)^{2}}\Big{[}&C_{J1,2}\dot{\varphi}_{1}^{2}+C_{2}\dot{\varphi}_{2}^{2}+C_{\textnormal{J3}}(\dot{\varphi}_{2}-\dot{\varphi}_{1})^{2}+C_{\textnormal{23}}(\dot{\varphi}_{3}-\dot{\varphi}_{2})^{2}+C_{3}\dot{\varphi}_{3}^{2}+C_{\textnormal{34}}(\dot{\varphi}_{4}-\dot{\varphi}_{3})^{2}+\\ &+C_{4}\dot{\varphi}_{4}^{2}+C_{5}\dot{\varphi}_{5}^{2}+C_{\textnormal{45}}(\dot{\varphi}_{5}-\dot{\varphi}_{4})^{2}+C_{\textnormal{35}}(\dot{\varphi}_{5}-\dot{\varphi}_{3})^{2}+C_{6}\dot{\varphi}_{6}^{2}+\\ &+C_{\textnormal{56}}(\dot{\varphi}_{6}-\dot{\varphi}_{5})^{2}+C_{\textnormal{26}}(\dot{\varphi}_{6}-\dot{\varphi}_{2})^{2}+C_{\textnormal{J8}}(\dot{\varphi}_{7}-\dot{\varphi}_{6})^{2}+C_{J9,10}\dot{\varphi}_{7}^{2}\Big{]},\end{split}$ (12)
@@
Figure 6: Coupler energy levels as a function of the external magnetic flux in the coupler SQUID and corresponding system states. The coupler state is determined by the states of three subsystems. The three numbers reflect the number of excitations in each of the coupler modes. Levels are split due to the capacitive interaction between the nodes.
Figure 7: (a) Energy levels of the transmon-coupler-transmon system from external magnetic flux in the coupler with parameters from Tab.II. The system is determined by the states of five subsystems. The first symbol in the system quantum state corresponds to the state of the first transmon ($|g\rangle$, $|e\rangle$ or $|f\rangle$), the second symbol corresponds to the state of the second transmon (similarly to the first transmon), the rest corresponds to the state of the coupler as in Fig. 6. Coupler tunable mode pushes transmon levels, changing the value of ZZ-interaction between qubits. The system states that are important when considering population dynamics (see Fig. 8) are noted. (b) ZZ-interaction between qubits vs external magnetic flux on the coupler. The red circle shows the minimal value of ZZ-interaction.
@@
where $C_{\text{new}}^{-1}$ is the inverse capacitance matrix. We calculate the eigenvalues of each subsystem in phase basis using numerical diagonalization and find charge matrix elements by means of Fourier transformation. Then we construct the full Hamiltonian introducing the interaction terms in the form of Eq. (2) from the main text. Charge energy here is defined as $E_{Cij}=\frac{e^{2}}{2}C_{\text{new}}^{-1}.$ Finally, we numerically diagonalized the full Hamiltonian.
@@
The calculated energy spectrum as a function of magnetic flux $\Phi$ in the coupler SQUID is plotted in Fig. 7(a). We provide the full system spectrum calculated for the qubits at the operating point as a function of an external magnetic flux $\Phi_{\text{ext}}$. The eigenstates are labelled as $|n_{\text{q1}},n_{\text{q2}},n_{\text{c1}},n_{\text{c2}},n_{\text{c3}}\rangle$, where $n_{i}$ is the occupancy of the $i$-th mode. The four computational levels are $|gg000\rangle$, $|eg000\rangle$, $|ge000\rangle$, $|ee000\rangle$.

### v2

@@ III.3 CZ gate concept
We realize the two-qubit CZ gate with a flux pulse applied to the coupler SQUID. One can switch on the interaction adiabatically increasing the external flux and, after waiting for the required time, return the coupler to the idle bias point. The flux pulse should be as short as possible to avoid decoherence during the gate while remaining adiabatic in order to prevent undesired transitions between the computational states and leakage out of the computational subspace. Therefore, in order to activate the interaction we use a smooth pulse $\Phi_{\text{ext}}(t)$ with a Hann function-shaped edges of 9 ns both in the simulation and the experiment, as it is shown in Fig. 3(d). Also, we define the gate duration as the length of the entire pulse, accounting for both edges and a plateau-shaped part.
@@
We described above in details the method of constructing the system model using two-tone spectroscopy data and measurement of the ZZ interaction strength at the zero-flux point. Using the obtained values, we construct the full-circuit Hamiltonian and calculate the time dynamics of the system during the CZ gate implementation. In this numerical analysis, we take into account the first four energy levels of each qubit and each mode of the coupler in order to reduce a computational time. Thus, the Hilbert space dimension of the full system in the simulation is $4^{5}$.
Figure 3: The CZ gate simulation for the experimentally obtained device parameters. (a) The error (red), the swap error $\epsilon_{\mathrm{swap}}$ between the computational states (dashed blue) and the high-energy levels leakage error $\epsilon_{\mathrm{leak}}$ (dashed-dotted green) are shown as functions of the CZ gate duration. (b) The corresponding amplitude of the external flux in the coupler’s SQUID. (c) For the gate duration of 43.3 ns (star in panel (a)) the time evolution of the conditional two-qubit phase $\theta$ (red), the swap error $\epsilon_{\mathrm{swap}}$ (dashed blue), and the leakage error $\epsilon_{\mathrm{leak}}$ (dashed-dotted green) are simulated. The estimated fidelity estimated is $F=0.9997$. (d) The coupler flux pulse with Hann-shaped edges are used both in simulation and experiment.
@@
We simulate the CZ gate performance for a range of the coupler’s SQUID flux pulse amplitude from $0.37\Phi_{0}$ to $0.5\Phi_{0}$. For each amplitude value we find a gate duration, such that the conditional phase accumulated by the computational state $|ee\rangle$ is $\theta=\pi$. In order to evaluate the performance of the resulting CZ operation, we use the standard expression for the two-qubit gate fidelity utilizing Pauli transfer matrix representation [56, 57], which we then convert to the depolarizing error $\epsilon$. To analyze the error contribution, we calculate the average leakage error out of the computational subspace [58]
$\epsilon_{\mathrm{leak}}=1-\frac{1}{4}\sum_{i,j=1}^{4}\left|u_{ij}\right|^{2},$ (4)
and the swap error, $\epsilon_{\mathrm{swap}}$, which quantifies undesired for a CZ gate state population exchange within the computational subspace
$\epsilon_{\mathrm{swap}}=\frac{1}{4}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{4}\left|u_{ij}\right|^{2}.$ (5)
Here, $u_{ij}$ denotes the matrix elements of the evolution operator $\hat{\mathcal{U}}$ projected onto the two-qubit computational subspace, with the coupler in the ground state.
The resulting errors are presented in Fig. 3(a), while the corresponding gate duration as a function of an external flux amplitude is given in Fig. 3(b). Here we observe a periodic behavior of the gate error condutioned by population oscallations within the computational subspace. At the first minimum of the fidelity curve we do not get high-fidelity operation due to the immense ZZ interaction.
@@
Withal, the second error minimum, marked with a star, shows the gate fidelity of 0.9997 with 43.3-ns duration. For this point, a time evolution of the mean target state population (dashed blue), a conditional phase (red) and a leakage error (dashed-dotted green) are depicted in Fig. 3(c). The target state population behavior mostly relates to the oscillations between states during the gate implementation. The population leaks to the second excited qubit states $|gf\rangle,|fg\rangle$ and to the coupler state $|010\rangle$. The detailed dynamic of each computational state can be found in Appendix E.
III.4 CZ gate implementation
We implemented the proposed CZ gate with a total pulse duration of 60 ns, including 9-ns edge segments, at an amplitude of $-0.28$ V. Since at this amplitude the experimental ZZ interaction strength is slightly larger than in the model, the corresponding gate evolution in the experiment is shorter. Therefore, we assume that the performed gate corresponds to a point near the third minimum of the error function in Fig. 3(a). In order to validate the two-qubit gate performance we start with the quantum process tomography experiment, the result of which is depicted in Fig. 4(c). The fidelity of the reconstructed CZ Pauli transfer matrix is $83.3\%$. We attribute the low fidelity to measurement errors, specifically low-fidelity multi-qubit readout. Therefore, we utilize cross-entropy benchmarking (XEB) [1] resilient against SPAM errors in order to characterize the performance of the implemented CZ operation.
For the XEB we execute the gate sequence shown in Fig. 4(a), where $C_{ij}$ are randomly chosen single-qubit Clifford gates. Here the first index corresponds to the qubit number and the second index refers to the serial number of the gate in the sequence. The single-qubit Clifford group is constructed of the $\pi/2$ rotations around X axis of $46.66$ ns duration and virtual S gates in the same way as it was performed in Ref. [59]. First, we accomplish reference XEB consisting of only single-qubit operations simultaneously executed on both qubits. The obtained depolarization fidelity averaged over randomly sampled quantum circuits as a function of a circuit depth $m$ is shown in Fig. 4(b) with a blue color. We approximate the data with the function $ap^{m}$, where $p$ is the depolarizing parameter and $a$ is a fitting parameter used to take into account a state preparation and measurement (SPAM) errors. The resulted depolarization parameter $p_{1}$ is $0.9714\pm 0.0005$. The similar data collected for the same circuits with the interleaved CZ gate is depicted in Fig. 4(b) in green. The depolarization parameter $p_{2}$ is equal to $0.947\pm 0.001$. In order to calculate the conventional fidelity $F$ of a target two-qubit gate we use the formula [1, 60]
@@ III.4 CZ gate implementation
where $p=p_{2}/p_{1}$. The obtained CZ gate fidelity $F_{\mathrm{CZ}}$ is $(98.14\pm 0.12)\%$. A detailed analysis of errors, along with the corresponding calculations, is provided in Appendix F. Below, we outline the key points.
The total error in interleaved benchmarking arises from single-qubit gate errors and two-qubit CZ gate errors. Single-qubit errors primarily caused by relaxation, dephasing, and static ZZ interaction processes. Based on measured relaxation and coherence times (Table 1), we estimate a depolarization error of 0.015 due to amplitude and phase damping, while the contribution from static ZZ interaction adds 0.012. The combined effect yields an overall single-qubit gate fidelity of approximately 98.0%, which agrees well with experiment.
For the CZ gate, error analysis is more complex due to the interplay of relaxation, dephasing, conditional phase error, leakage, and state population exchange. We begin by estimating the error arising from qubit relaxation and dephasing. In preliminary measurements, we observe that the coherence time $T_{2}^{*}$ decreases by approximately a factor of two when the coupler is tuned to the gate activation flux point, while the relaxation time remains stable across the operating flux range. Assuming that during gate operation $T_{2}^{*}$ is reduced by half compared to its value in the idle state, we estimate a depolarization error of 0.016. The phase error $\delta\theta=0.03\pi$, estimated by XEB method, contributes an additional error of less than 0.002. Leakage out of the computational subspace is assessed by extracting state purity from XEB data using the speckle purity benchmarking technique, yielding a leakage error of 0.008.
Taken together, these results highlight decoherence processes as the dominant error source and particularly underscore the need for suppressing the static ZZ interaction.
Qubit $f_{q}$ $\delta_{q}$ $T_{2}^{*}$ $T_{1}$ $t_{\text{1Q}}$ $F_{\text{1Q}}$
(GHz) (MHz) $(\mu s)$ $(\mu s)$ (ns) (%)
A 6.450 -156 8.6 15.8 $2\cdot 46.66$ 99.1
B 6.494 -157 10.8 8.3 $2\cdot 46.66$ 98.3
Table 1: The device parameters: qubits frequency, anharmonicity, dephasing and coherence times, and the length of single-qubit Clifford gates in Fig. 4(a). The coefficient 2 denotes the fact, that we construct the Clifford group out of two $R_{\mathrm{X}}(\pi/2)$ gates with duration of 46.66-ns. The single-qubit gate fidelities are obtained by cross-entropy benchmarking performed simultaneously on both qubits.
@@ Acknowledgments
The authors acknowledge Alexey Ustinov for fruitful discussions and useful comments on the manuscript. The work was supported by Rosatom in the framework of the Roadmap for Quantum computing (Contract No. 151/21-503, December 21, 2021 dated 12/21/2021). Theoretical analysis is partially supported by the Priority 2030 program at the National University of Science and Technology “MISIS” (Strategic Project Quantum Internet). The sample was fabricated using the equipment of MIPT Shared Facilities Center (Contract No.868/221-D dated October 24, 2022).
@@ Appendix A Full-circuit Hamiltonian
Josephson Energy (GHz) Josephson Capacitance (fF) Node Capacitance (fF)
Design Experiment Design Experiment
$E_{J1}$ 124.2 104.4 $C_{J1}$ 11.1 9.3 $C_{2}$ 75.6
$E_{J2}$ 43.5 23.1 $C_{J2}$ 3.9 2.1 $C_{3}$ 47.2
$E_{J3}$ 17.4 28.8 $C_{J3}$ 1.5 2.6 $C_{4}$ 144
$E_{J4}$ 124.2 159.8 $C_{J4}$ 11.1 14.2 $C_{5}$ 46.5
$E_{J5}$ 79.5 191.5 $C_{J5}$ 11.5 17.1 $C_{6}$ 75.9
$E_{J6}$ 129.1 117.0 $C_{J6}$ 7.1 10.4 $C_{23}$ 9
$E_{J7}$ 124.2 159.8 $C_{J7}$ 11.1 14.2 $C_{26}$ 0.015
$E_{J8}$ 17.4 29.3 $C_{J8}$ 1.5 2.6 $C_{34}$ 13
$E_{J9}$ 43.5 23.5 $C_{J9}$ 3.9 2.1 $C_{35}$ 0.005
$E_{J10}$ 124.2 106.1 $C_{J10}$ 11.1 9.5 $C_{56}$ 9.5
Table 2: Target circuit parameters of the device: Josephson energies of junctions, node and Josephson junction capacitances. The design parameters of Josephson junctions are compared with the values extracted from the experiment. The coupler Josephson junction capacitances $C_{J4}$, $C_{J5}$, $C_{J6}$, and $C_{J7}$ are included in corresponding values $C_{34}$, $C_{4}$, and $C_{45}$.
@@
The single-qubit Hamiltonian has the following form [47]:
@@
$\begin{split}T=\frac{\Phi^{2}_{0}}{2(2\pi)^{2}}\Big{[}&C_{\textnormal{J1,2}}\dot{\varphi}_{1}^{2}+C_{2}\dot{\varphi}_{2}^{2}+C_{\textnormal{J3}}(\dot{\varphi}_{2}-\dot{\varphi}_{1})^{2}+C_{\textnormal{23}}(\dot{\varphi}_{3}-\dot{\varphi}_{2})^{2}+C_{3}\dot{\varphi}_{3}^{2}+C_{\textnormal{34}}(\dot{\varphi}_{4}-\dot{\varphi}_{3})^{2}+\\ &+C_{4}\dot{\varphi}_{4}^{2}+C_{5}\dot{\varphi}_{5}^{2}+C_{\textnormal{45}}(\dot{\varphi}_{5}-\dot{\varphi}_{4})^{2}+C_{\textnormal{35}}(\dot{\varphi}_{5}-\dot{\varphi}_{3})^{2}+C_{6}\dot{\varphi}_{6}^{2}+\\ &+C_{\textnormal{56}}(\dot{\varphi}_{6}-\dot{\varphi}_{5})^{2}+C_{\textnormal{26}}(\dot{\varphi}_{6}-\dot{\varphi}_{2})^{2}+C_{\textnormal{J8}}(\dot{\varphi}_{7}-\dot{\varphi}_{6})^{2}+C_{\textnormal{J9,10}}\dot{\varphi}_{7}^{2}\Big{]},\end{split}$ (13)
@@
Figure 6: (a) Energy levels of the transmon-coupler-transmon system from external magnetic flux in the coupler with parameters from Tab.II. The system is determined by the states of five subsystems. The first symbol in the system quantum state corresponds to the state of the first transmon ($|g\rangle$, $|e\rangle$ or $|f\rangle$), the second symbol corresponds to the state of the second transmon (similarly to the first transmon), the rest corresponds to the state of the coupler. Coupler tunable mode pushes transmon levels, changing the value of ZZ interaction between qubits. The system states that are important when considering population dynamics (see Fig. 9) are noted. (b) ZZ interaction between qubits vs external magnetic flux on the coupler. The inset shows the coupling near the zero flux bias point, illustrating the rate of its increase. The green circle shows the minimal value of ZZ interaction.
@@
where $C_{\text{new}}^{-1}$ is the inverse capacitance matrix. We calculate the eigenvalues of each subsystem in phase basis using numerical diagonalization and find charge matrix elements by means of Fourier transformation. Then we construct the full Hamiltonian
$\hat{H}_{\text{int}}=\sum_{i\neq j}E_{Cij}\hat{n}_{i}\hat{n}_{j},$ (22)
where $\{i,j\}$ is the set of circuit nodes pairs, $\hat{n}$ is the Cooper pair number operator and $E_{Cij}$ is the energy of the capacitive interaction. Charge energy here is defined as $E_{Cij}=\frac{e^{2}}{2}C_{\text{new}}^{-1}.$
@@
Finally, we numerically diagonalized the full Hamiltonian. The calculated energy spectrum as a function of magnetic flux $\Phi$ in the coupler SQUID is plotted in Fig. 6(a). We provide the full system spectrum calculated for the qubits at the operating point as a function of an external magnetic flux $\Phi_{\text{ext}}$. The eigenstates are labelled as $|n_{\text{q1}},n_{\text{q2}},n_{\text{c1}},n_{\text{c2}},n_{\text{c3}}\rangle$, where $n_{i}$ is the occupancy of the $i$-th mode. The four computational levels are $|gg000\rangle$, $|eg000\rangle$, $|ge000\rangle$, $|ee000\rangle$.